OpenStudy (anonymous):

help..question uploaded..AP CALCULUS QUIZ QUESTIONS both images go to the same information that is provided above 12.

OpenStudy (istim):

Hey, I'm doing Calculus too. We should learn from each other.

OpenStudy (anonymous):

lol i hardly understand ap

OpenStudy (istim):

Well, with the help from each other, we can go form being mediocre to decent, and then up from there.

OpenStudy (anonymous):

right now i just want to figure this out to be honest

OpenStudy (apoorvk):

dy/dx = (1/2) ycosx dy/y = (cosx.dx)/2 Integrate both sides. and get lny = sinx/2 + c (where is c is constant of integration) Now, since its given f(0)=2, plug in x=0 and y=2 in the solved equation above to find out what 'c' is. so, ln2 = c So, the differential equation is: lny = sinx/2 +ln2

OpenStudy (apoorvk):

*solved differential equation.

OpenStudy (anonymous):

thats not even an answer choice

OpenStudy (istim):

Maybe...we can derive an answer from that?

OpenStudy (anonymous):

if he is doing 13 i got d.

OpenStudy (anonymous):

omg this is hard.

OpenStudy (apoorvk):

Oh sorry, the image was quite big, so I didn't read the whole question, thinking that was all, (I scrolled down later). I 'll post a solution for that, once the website restarts.

OpenStudy (anonymous):

alright thank you! :) sorry i did not know the pic was big

OpenStudy (apoorvk):

Its okay, it was my foolishness not to have scrolled down :P